### v1

# Direct Unlearning Optimization for Robust and Safe Text-to-Image Models

Abstract
Recent advancements in text-to-image (T2I) models have greatly benefited from large-scale datasets, but they also pose significant risks due to the potential generation of unsafe content. To mitigate this issue, researchers have developed unlearning techniques to remove the model’s ability to generate potentially harmful content. However, these methods are easily bypassed by adversarial attacks, making them unreliable for ensuring the safety of generated images. In this paper, we propose Direct Unlearning Optimization (DUO), a novel framework for removing Not Safe For Work (NSFW) content from T2I models while preserving their performance on unrelated topics. DUO employs a preference optimization approach using curated paired image data, ensuring that the model learns to remove unsafe visual concepts while retaining unrelated features. Furthermore, we introduce an output-preserving regularization term to maintain the model’s generative capabilities on safe content. Extensive experiments demonstrate that DUO can robustly defend against various state-of-the-art red teaming methods without significant performance degradation on unrelated topics, as measured by FID and CLIP scores. Our work contributes to the development of safer and more reliable T2I models, paving the way for their responsible deployment in both closed-source and open-source scenarios.
CAUTION: This paper includes model-generated content that may contain offensive or distressing material.
{NoHyper} Work done during an internship at Naver AI Lab {NoHyper} Corresponding authors
1 Introduction
In recent years, text-to-image (T2I) models [21, 46, 44, 39, 8, 7] have experienced significant advancements thanks to large-scale data and diffusion models. However, the large-scale web-crawled data, e.g., LAION [50] often include a significant amount of inappropriate and objectionable material, so T2I models trained on such data may pose a potential risk of generating harmful content such as Not Safe For Work (NSFW) content, copyright infringement, and a violation of privacy [45, 4]. Filtering and curating such large-scale training data [54] is not feasible for large-scale datasets. At the production level, service providers usually block inappropriately generated images with ad-hoc classifier [54], but attackers can easily bypass the classifier by accessing the model directly.
To address the generation of unsafe content, there has been a substantial amount of research focusing on directly fine-tuning diffusion models to forget unsafe concepts [11, 12, 25, 32, 16, 35, 62]. Recent work in the field has primarily focused on prompt-based approaches to unlearning, which aim to induce models to behave as if they had been provided with safe prompts when an unsafe prompt is given [11, 25, 32, 62, 35, 12]. These methods effectively censor NSFW image generation without degrading the image generation capability. However, recent studies [41, 55, 59] warn that by introducing slight perturbations to the unsafe prompts, it is possible to circumvent these censoring mechanisms and generate undesirable images (Figure 1).
Figure 1: Image-based unlearning. Prompt-based unlearning can be easily circumvented with adversarial prompt attack. On the other hand, image-based unlearning robustly produces safe images regardless of the given prompt. We use Refer to caption for publication purposes.
Current unlearning methods are vulnerable to adversarial attacks because it is difficult to cover all NSFW concepts using prompts alone. For instance, even if the prompt “nudity” is removed, many related concepts like “naked,” “erotic,” or “sexual” may remain. Furthermore, it is infeasible to anticipate and eliminate all possible adversarial prompts beforehand. Moreover, unlearning methods that rely on text conditioning cause the model to behave as if it cannot comprehend the unsafe prompts, rather than removing the internal visual concepts within the model.
To fundamentally address limitations, it is necessary to guide the model to remove unsafe visual features regardless of the prompt. One way to achieve this is by directly unlearning the unsafe images instead of prompts. However, simply moving away from the target image leads to ambiguity regarding which concepts to forget and can cause forgetting of unrelated visual features. For instance, if we make the model unlearn the image in Figure 3-(a), it would lose the ability to generate not only nudity but also unrelated concepts like women or forests.
To mitigate this ambiguity, we reformulate the unlearning as a preference optimization problem. Preference optimization is a technique aimed at training models to generate desirable outputs when provided with both positive and negative examples. Our main intuition is that, unlike when only an undesirable sample is given, the presence of a desirable sample helps the model resolve the ambiguity about what information should be retained from the undesirable sample. As shown in Figure 3-(b), thanks to the desirable image, the model can recognize that the forest and woman are not subjects that need to be erased.
In this paper, we propose Direct Unlearning Optimization (DUO), a method to remove unsafe visual features directly from the model while maintaining generation quality for unrelated topics. Specifically, we created paired data by using unsafe images and their counterparts, from which unsafe concepts were removed using the SDEdit [33]. Then, we employ Direct Preference Optimization (DPO) [43] to make the model favor the latter group while encouraging it to move away from the former group. As shown in Figure 3-(b), preference optimization through a paired dataset explicitly specifies which features of a given image should be forgotten. To further regularize the model not to unlearn for unrelated topics, we introduce output-preserving regularization [36, 61, 6] that forces to preserve the denoising capability of the model. We call our framework as Direct Unlearning Optimization (DUO). We demonstrate that our method can robustly defend against various state-of-the-art red teaming methods [55, 41] without significant performance degradation in unrelated topics, as measured by LPIPS [64], FID [18] and CLIP scores [42]. Additionally, through an ablation study, we validate the efficacy of output preservation regularization and visual feature unlearning.
(a) Unsafe images contain both unsafe concepts and unrelated concepts.
(b) Excluding unrelated concepts by utilizing both unsafe images and their corresponding safe images.
Figure 3: Importance of using both unsafe and paired safe images to preserve model prior. We use Refer to caption for publication purposes. Unsafe concept refers to what should be removed from the image (red), while unrelated concept refers to what should be retained in the image (green).
2 Related work
Text-to-Image (T2I) models with Safety Mechanisms.
The safety mechanisms designed to prevent undesirable behavior in T2I models can be classified into four categories: dataset filtering [54, 14], concept unlearning [11, 12, 25, 32, 16, 35, 62], in-generation guidance [49], and post-generation content screening [1, 54, 3]. These methods are orthogonal to each other, and using them together can make the T2I generation system safer.
Due to the advantage of not requiring training from scratch, fine-tuning-based unlearning has been actively researched recently. Pioneering works like ESD [11] and CA [25] train the model to generate the same noise regardless of whether an unsafe prompt is present or not. SPM [32] trains a one-dimensional regulator instead of full model training. TIME [35], UCE [12] and MACE [30] update prompt-dependent cross-attention model weights with closed-form equation. Forget-me-not [62] suppresses the attention map of cross-attention layers for unsafe prompts. SA [16] proposed an unlearning method based on continual learning, utilizing Elastic Weight Consolidation (EWC). These works rely on prompts during training, making them vulnerable to adversarial prompt attacks, as discussed later.
Meanwhile, SafeGen [28], as a concurrent work with ours, addresses similar concerns regarding adversarial prompts and utilizing only images for training. They use blurred images and guide a model using supervised learning. In contrast to their work, we utilize paired data generated with SDEdit [33] to enable the model to erase only unsafe concepts and employ preference optimization. Therefore, we have the advantage of not requiring separate benign data and producing erased results more closely resemble real images.
Red-Teaming for T2I models.
Red teaming is a method for searching for vulnerabilities in security systems. Extending this concept to machine learning, many red teaming techniques have been proposed to explore how robust a model’s safety mechanism is. For Text-to-Image (T2I) models, studies have shown that even with safety mechanisms applied, it is still possible to generate harmful images through prompt engineering [55, 41, 59, 9, 66, 15]. Ring-A-bell [55] and SneakyPrompt [59] generate adversarial prompts in a black box scenario without access to the denoising model. They collect unsafe prompts and find prompts with similar embeddings through optimization. On the other hand, Concept Inversion [41] is a method that can be used in a white box scenario where access to the gradient of the denoising model is available. This method learns a special token $\langle c\rangle$ through textual inversion that contains information about unsafe images, and then uses it to generate unsafe images. The goal of our research is to propose a safety mechanism that is robust against red teaming.
Preference optimization in T2I models.
Direct Preference Optimization (DPO) [43] has been frequently used in language models as it enables preference-based model tuning without the need for reward models. DiffusionDPO [56] and D3PO [58] extend this approach to diffusion models, ensuring that generated images better reflect user preferences. Diffusion-KTO [27] replaces DPO with Kahneman-Tversky Optimization (KTO) [10], eliminating the need for paired data. DCO [26] applies preference optimization to the personalization domain, ensuring that models preserve priors while generating personal images. Our work is the first to apply preference optimization to the unlearning problem and proposes task-specific data and regularization methods to fit unlearning task.
3 Method
In this section, we introduce our framework named Direct Unlearning Optimization (DUO) consisting of Direct Preference Optimization (DPO) with synthetic paired data and output-preserving regularization. After providing a preliminary explanation of diffusion models, we detail the process of generating paired images, which are necessary for successful preference optimization for concept unlearning. Next, we provide a detailed explanation of applying direct preference optimization to the unlearning task using the generated paired images. Finally, we introduce output-preserving regularization loss, a novel strategy to maintain the prior distribution of the model.
3.1 Preliminary: Diffusion models
Diffusion models [51, 20, 52, 53, 31] are generative models that learn to generate images through an iterative denoising process. For an image $x_{0}$, the diffusion process samples a noisy image $x_{t}$ from $q(x_{t}|x_{0})=\mathcal{N}(\sqrt{\alpha_{t}}x_{0},\sqrt{1-\alpha_{t}}I)$. The magnitude of the added noise is determined by the noise schedule $\alpha_{t}$, and it increases as $t$ becomes larger. We train the model $\epsilon_{\theta}(\cdot)$ to denoise the noisy image $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon$ by predicting the added noise $\epsilon$.
$L_{\text{DSM}}=\mathbb{E}_{x_{0}\sim q(x_{0}),x_{t}\sim q(x_{t}|x_{0})}[||\epsilon-\epsilon_{\theta}(x_{t})||_{2}^{2}]$ (1)
This is called the denoising score matching (DSM) loss. It encourages the model to estimate the gradient of the log-probability of the noisy data, i.e., $\epsilon_{\theta}(x_{t})\propto\nabla\log p(x_{t}\mid c)$.
Diffusion models can be conditioned on other modalities, such as text, to generate images that align with the given context [19, 34, 63]. In this case, the model learns to estimate $\epsilon(x_{t},c)\propto\nabla\log p(x_{t})$, where $c$ represents the conditioning information. This allows the model to generate images that are not only realistic but also semantically consistent with the provided text or other modalities.
3.2 Synthesizing paired image data to resolve ambiguity in image-based unlearning
(a) Original unsafe image
(b) Prompt substitution
(c) SDEdit method
Figure 5: Effectiveness of utilizing SDEdit for generating paired image data for unlearning. When unlearning (a), if a retain sample is naively generated as shown in (b), the model cannot accurately determine what information should be retained and what should be forgotten. Using SDEdit as shown in (c), the model can identify what information should be retained and what should be forgotten from the undesirable sample. We use Refer to caption for publication purposes.
As we mentioned in Section 1, even images with unsafe concepts contain numerous unrelated visual features that need to be preserved (e.g., photo-realism, background, and humans). Therefore, a method is needed to selectively remove only the visual features associated with the targeted unsafe concept. Intuitively, if we provide the model with images that have identical unrelated visual features but differ in the targeted unsafe concept, the model can be guided to unlearn only the features related to the unsafe concept. Unfortunately, as shown in Figure 5 (b), naively generating paired images with different prompts often resulted in changes to the unrelated attributes of the images.
We solve this by creating image pairs using SDEdit [33], a method for image translation. The process of creating image pairs with SDEdit is as follows. Given an undesirable concept $c^{-}$, e.g., naked, we first generate an unsafe image $x_{0}^{-}$ using $c^{-}$. Then, we add a certain amount of Gaussian noise to send it to $x_{t}^{-}$. Lastly, denoise $x_{t}^{-}$ to create paired safe image $x_{0}^{+}$ with negative guidance for $c^{-}$. If possible, we also use positive guidance for $c^{+}$, e.g., dressed, during the denoising process. Thanks to the characteristics of diffusion models, SDEdit allows us to generate images that have similar coarse features but either include or exclude the unsafe concept, ensuring that the concept unlearning process is focused on the specific elements we want to remove while preserving the rest of the image as shown in Figure 5 (c).
Now, we want to make model not to generate visual features in unsafe image $x_{0}^{-}$ while still can generate safe image $x_{0}^{+}$. We solve this as a preference optimization problem. In the next subsection, we will discuss how we consider our generated paired dataset as preference data and apply Direct Preference Optimization (DPO) to our unlearning problem.
3.3 Concept unlearning as a preference optimization problem
In subsection 3.2, we ensure that the paired images only differed in the presence or absence of unsafe concepts while keeping other attributes the same. Using them, we need to guide the model to remove only the visual features associated with unsafe concepts. We employed a preference optimization method where the model is trained to move away from unsafe image $x_{0}^{-}$ and towards their corresponding safe images $x_{0}^{+}$. Since the only difference between our dataset pairs $x_{0}^{-}$ and $x_{0}^{+}$ is the information related to the unsafe concept, the unrelated concepts remain unaffected.
Suppose we are given a paired dataset $\{x_{0}^{+},x_{0}^{-}\}$, where $x_{0}^{+}$ and $x_{0}^{-}$ denote the retain and forget sample, respectively. The probability of preferring $x_{0}^{+}$ over $x_{0}^{-}$ can be modeled by the reward function $r(x)$:
$p(x_{0}^{+}\succ x_{0}^{-})=\sigma(r(x_{0}^{+})-r(x_{0}^{-}))$ (2)
where $\sigma$ is the sigmoid function. This model is known as the Bradley-Terry (BT) model [5]. The reward function can be learned using a binary cross-entropy loss to maximize likelihood:
$L_{\text{BT}}(r)=-\mathbb{E}_{x_{0}^{+},x_{0}^{-}}[\log\sigma(r(x_{0}^{+})-r(x_{0}^{-}))]$ (3)
Given this reward function, preference optimization aims to fine-tune the model to maximize the reward. To prevent the model $p_{\theta}(x)$ from excessively forgetting its existing capabilities, we use KL-constrained reward optimization as our objective:
$\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}[r(x_{0})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0})||p_{\phi}(x_{0})]$ (4)
where $p_{\phi}$ denotes the pretrained model distribution, and the hyperparameter $\beta$ controls the extent to which the model diverges from the prior distribution. The above equation has a unique closed-form solution $p_{\theta}^{*}$ [13, 24, 38, 40]:
$p_{\theta}^{*}=p_{\phi}(x_{0})\exp(r(x_{0})/\beta)/Z$ (5)
where $Z=\sum_{x_{0}}p_{\phi}(x_{0})\exp(r(x_{0})/\beta)$ is the partition function. By manipulating this equation, we can express the reward function in terms of the fine-tuned and pretrained model distributions:
$r(x_{0})=\beta\log{\frac{p_{\theta}^{*}(x_{0})}{p_{\phi}(x_{0})}}+\beta\log Z$ (6)
Plugging this into the BT loss equation, we obtain the DPO loss [43]:
$L_{\text{DPO}}=-\mathbb{E}_{x_{0}^{+},x_{0}^{-}}[\log\sigma(\beta\log{\frac{p_{\theta}^{*}(x_{0}^{+})}{p_{\phi}(x_{0}^{+})}}-\beta\log{\frac{p_{\theta}^{*}(x_{0}^{-})}{p_{\phi}(x_{0}^{-})}})]$ (7)
Diffusion-DPO.
[…]
$r(x_{0})=\mathbb{E}_{p_{\theta}(x_{1:T}|x_{0})}[R(x_{0:T})]$ (8)
[…]
$\max_{p_{\theta}}\mathbb{E}_{x_{0:T}\sim p_{\theta}(x_{0:T})}[R(x_{0:T})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T})||p_{\phi}(x_{0:T})]$ (9)
Following a similar process as with the DPO objective, we can derive Diffusion-DPO objective to optimize $p_{\theta}(x_{0:T})$:
$L_{\text{Diffusion-DPO}}=-\mathbb{E}_{x_{0}^{+},x_{0}^{-}}[\log\sigma{(}\mathbb{E}_{p_{\theta}(x_{1:T}^{+}|x_{0}^{+}),p_{\theta}(x_{1:T}^{-}|x_{0}^{-})}[\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{+})}{p_{\phi}(x_{0:T}^{+})}}-\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{-})}{p_{\phi}(x_{0:T}^{-})}}]{)}]$ (10)
By leveraging Jensen’s inequality and the ELBO, this loss can be expressed as a combination of fully tractable denoising score matching losses (see the Appendix C for detailed derivation):
$L_{\text{Diffusion-DPO}}\leq$ $-\mathbb{E}_{(x_{t}^{+},x_{t}^{-})\sim D,x_{t}^{+}\sim q(x_{t}^{+}|x_{0}^{+}),x_{0}^{-}\sim q(x_{t}^{-}|x_{0}^{-})}$ (11)
$\left[\log\sigma\left(-\beta\left(\begin{aligned} &\|\epsilon-\epsilon_{\theta}(x_{t}^{+},t)\|_{2}^{2}-\|\epsilon-\epsilon_{\phi}(x_{t}^{+},t)\|_{2}^{2}\\ &-\left(\|\epsilon-\epsilon_{\theta}(x_{t}^{-},t)\|_{2}^{2}+\|\epsilon-\epsilon_{\phi}(x_{t}^{-},t)\|_{2}^{2}\right)\end{aligned}\right)\right)\right]$
In Eq. (17), $\|\epsilon-\epsilon_{\theta}(x_{t}^{+},t)\|_{2}^{2}-\|\epsilon-\epsilon_{\phi}(x_{t}^{+},t)\|_{2}^{2}$ operates as gradient descent for the preferred sample, while $\|\epsilon-\epsilon_{\theta}(x_{t}^{-},t)\|_{2}^{2}-\|\epsilon-\epsilon_{\phi}(x_{t}^{-},t)\|_{2}^{2}$ operates as gradient ascent for the dispreferred sample.
3.4 Output-preserving regularization
Unfortunately, in preliminary experiments, we discovered that DPO’s KL divergence regularization alone is insufficient for prior preservation. The main reason was that the gradient ascent term in Eq. (17) does not simply unlearn $x_{0}^{+}$, but also decreases the model’s denoising ability.
To mitigate this issue, we introduced the following novel regularization term to maintain the diffusion model’s denoising capability:
$L_{\text{prior}}=||\epsilon_{\phi}(x_{T})-\epsilon_{\theta}(x_{T})||_{2}^{2}$ (12)
The above regularization maintains the output even when unlearning is performed and the image is completely noise, i.e., at $x_{T}$. The reason for limiting it to $t=T$ is that when image information remains, output preservation can interfere with unlearning. On the other hand, when $t=T$, only complete noise information remains in the image, so we expected that enforcing output preservation would not affect unlearning.
Our final loss function, which incorporates the output preservation regularization, is as follows:
$L_{\text{DUO}}\triangleq$ $-\mathbb{E}_{x_{t}^{+}\sim q(x_{t}^{+}|x_{0}^{+}),x_{0}^{-}\sim q(x_{t}^{-}|x_{0}^{-}),x_{T}\sim\mathcal{N}(0,I)}$ (13)
$\left[\log\sigma\left(-\beta\left(\begin{aligned} &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\epsilon-\epsilon_{\theta}(x_{t}^{+},t)\|_{2}^{2}-\|\epsilon-\epsilon_{\phi}(x_{t}^{+},t)\|_{2}^{2}}\\ &-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\|\epsilon-\epsilon_{\theta}(x_{t}^{-},t)\|_{2}^{2}+\|\epsilon-\epsilon_{\phi}(x_{t}^{-},t)\|_{2}^{2}\right)}\end{aligned}\right)\right)+\lambda L_{\text{prior}}\right]$
4 Experiments
4.1 Experiments setup
Unlearning setup.
We use Stable Diffuson 1.4v (SD1.4v) with a LoRA [22, 47] with the Adam optimizer for fine-tuning. [11] We do not train the cross-attention layers to minimize the dependency on the text prompt for unlearning. Our experiments are conducted on Stable Diffusion version 1.4 to compare with other baseline methods. We use ESD [11], UCE [12], SPM [32] and MACE [30] as baselines for the concept unlearning category of the safety mechanism. SPM has a Facilitated Transfer (FT) mechanism that adjusts the strength of unlearning based on the similarity of given prompt to the unlearned concepts. This mechanism can be disadvantageous against adversarial prompt attacks, so we additionally report results from SPM without this mechanism (SPM w/o FT). For implementation details, please refer to the Appendix.
Red teaming.
For black box red teaming methods that do not access model weights, we use Ring-A-Bell [55] and SneakyPrompt [59]. Ring-A-Bell uses a genetic algorithm to find prompts that have similar embeddings to unsafe prompts defined by the attacker. We used 95 pre-trained prompts for nudity and 250 prompts for violence provided by the authors. SneakyPrompt slightly perturbs undesirable prompts that are blocked by the safety filter, e.g., "A naked man riding a bike.", to bypass the safety filter, e.g., "A grponypui man riding a bike". We used 200 prompts trained by the authors on SD1.4v for evaluating SneakyPrompt on nudity.
For white-box attacks that access model weights, we use Concept Inversion [41]. The concept inversion protocol is as follows. First, malicious images are generated using prompts provided by the i2p benchmark [37]. For nude images, prompts in the sexual category are used, and for violence, prompts with a toxicity probability of 0.95 or higher are used. For the images obtained through this process, a special token <c> is trained using textual inversion with a model that has a safety mechanism applied. The trained special token is then used as a prefix for the sexual/toxic prompts from the i2p benchmark.
Evaluation.
To evaluate whether we have defended against red teaming, we use external models. If these models determine that the image does not contain the unsafe concept, we consider the defense successful and report this ratio as the defense success rate. To classify whether images contain nudity, we use the NudeNet detector [2]. Following Tsai et al. [55], if the detector detects one of the following labels in the image: "FEMALE BREAST EXPOSED, FEMALE GENITALIA EXPOSED, ANUS EXPOSED, MALE GENITALIA EXPOSED", we consider the image to contain nudity.
To classify whether images contain violence, we used GPT-4o, a state-of-the-art multimodal LLM [57, 60, 1]. We use the user prompt, "Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? Please answer yes or no." [48] along with the image. If the model’s response is “yes”, we determine that it contains violence, while “no” indicated unlearning success. If neither token is present, we regenerate the model output. For more discussion on violence evaluation, please refer to the appendix (WIP).
To evaluate model performance unrelated to unlearned concept, we measure the FID [18] and CLIP scores [17] using MS COCO 30k validation dataset [29]. These metric provide insights into the capabilities of unlearned models. Additionally, we report the LPIPS [64] between images from the original SD1.4v model with those from the unlearned model using identical noise. This quantifies how the unlearned model’s distribution differs from the original.
4.2 Red teaming results
Figure 6: Quantitative result on nudity. The defense success rate (DSR) refers to the proportion of desirable concepts are generated. Prior preservation represents 1 - LPIPS between images generated by the prior model and the unlearned model. Results closer to the top right indicate better outcomes.
(a) Ring-A-Bell
(b) Concept Inversion
Figure 8: Qualitative result on nudity. Ours We use Refer to caption for publication purposes.
Unlearning is a task to satisfy the trade-off between unlearning performance and prior preservation simultaneously. Therefore, we report the Pareto curve according to the KL regularization. We plot the Pareto curve by adjusting the magnitude of KL regularization. Specifically, each point in the figures corresponds to $\beta\in\{500,250,100,50\}$.
Nudity.
Figure 6 shows the quantitative results for nudity detection. Our method maintains model performance comparable to the current state-of-the-art SPM in terms of prior preservation while achieving a defense success rate (DSR) of nearly 90% against all red teaming methods. A notable characteristic of the Pareto curve is the sharp bend where the increase in unlearning performance diminishes and the model rapidly loses its prior knowledge. Based on this observation, we recommend choosing $\beta=100$.
Figure 8 presents the qualitative results from Ring-A-Bell and concept inversion attacks. While previous prompt-based unlearning methods are easily circumvented with red teaming, our method demonstrates robust defense against these attacks. This result highlights the effectiveness of our image-based unlearning approach, which is independent of the prompt and robust against prompt-based red teaming. Interestingly, in the case of concept inversion, completely corrupted images are generated, unlike in the Ring-A-Bell attack. Through experimentation, we observed that it is challenging to robustly block textual inversion unless the generated result deviates significantly from the data manifold. This suggests that to defend against white box red teaming, which leverages gradient information, the internal features of the model must be completely removed rather than merely bypassed. Analyzing the different mechanisms of black box and white box red teaming presents an interesting direction for future research.
In Table 2, we report the FID and CLIP scores for MS COCO 30k. These scores indicate how well the unlearned model preserves its ability to generate images for unrelated prompts. Unlike other methods, DUO does not modify text embeddings or prompt-dependent weights; instead, it removes visual features directly from the model, making prior preservation quite challenging. Nonetheless, in terms of FID and CLIP scores, DUO shows performance comparable to other unlearning methods. Thus, for nudity, DUO is robust against prompt-based adversarial attacks while also preserving the original model’s generation capabilities for unrelated concepts. Figure 9 qualitatively confirms that the generation results do not significantly differ between the prior model and the model after unlearning.
Figure 9: Quanlitative result on prior preservation. The top row shows the original model, while the bottom row displays the results generated using prompts from the MS COCO validation 30k dataset after removing nudity. The same column uses the same initial noise and the same prompt.
Table 1: FID and CLIP score (CS) for nudity.
Method FID ($\downarrow$) CS ($\uparrow$)
ESD 14.05 29.97
UCE 14.30 30.64
SPM 13.74 30.92
SPM [^2] SPM without facilitate transfer 14.15 30.80
MACE 13.51 29.06
DUO 14.16 30.84
Table 2: FID and CLIP score (CS) for violence.
Method FID ($\downarrow$) CS ($\uparrow$)
ESD 21.33 27.35
UCE 16.20 30.40
SPM 13.60 30.89
SPM 13.44 30.84
MACE 14.85 29.14
DUO 17.52 30.34
Figure 10: Quantitative result on violence. The defense success rate (DSR) refers to the proportion of desirable concepts are generated. Prior preservation represents 1 - LPIPS between images generated by the prior model and the unlearned model. Results closer to the top right indicate better outcomes.
(a) Ring-A-Bell
(b) Concept Inversion
Figure 12: Qualitative result on violence. We used $\beta=500$ for Ring-A-bell and $\beta=1000$ for Concept Inversion. We use blurring for publication purposes.
Violence.
Figure 10 demonstrates the red teaming result on violence. Similar to nudity, DUO shows remarkable robustness against attacks while maintaining the same prior preservation score. We observe that the defense success rate values are generally not completely blocked compared to nudity. We believe is due to the fact that violence is a more abstract and complex concept compared to nudity, making unlearning more difficult. Figure 12 show qualitative results. We present the outcomes using $\beta=100$. Table 2 shows the FID and CLIP scores for MS COCO 30k. Similar to the nudity concept removal experiment, our method maintains comparable prior preservation performance to existing methods. Ulike the nudity experiment, in the violence concept removal experiment, we observe that the trade-off between prior preservation and defense success rate varies more steeply with the parameter $\beta$.
Figure 13: Ablation study on output preserving regularization (Ring-A-Bell). Output preserving regularization helps preserve the prior without significantly reducing the defense success rate.
4.3 Ablation study
Output preservation regularization.
We study the effect of output preservation regularization. As shown in Figure 13, lambda significantly enhances prior preservation performance at a similar level of DSR. This effect is more dramatic in the small $\beta$ regime. Without $L_{\text{prior}}$, the COCO generation result is severely degraded to the point of being unrecognizable. On the other hand, the unlearning result using $L_{\text{prior}}$ still produces plausible images at the same $\beta$. This demonstrates that output preservation regularization is effective in unlearning, where only the desired concept is removed. To assess the effectiveness of choosing the DPO method, we conducted an ablation study to evaluate the impact of KL divergence regularization. Please refer to Appendix A for ablation study for KL divergence regularization.
Visual feature unlearning.
The Figure 14 demonstrates the results when removing nudity using only the image compared to using both the image and a prompt. The red teaming was achieved through concept inversion. It is observed that when conducting unlearning using the prompt, DSR does not exceed 0.6. This shows vulnerability to adversarial prompts when utilizing prompt information.
5 Conclusion and limitation
Figure 14: Ablation study on visual feature unlearning (Concept Inversion). Using prompts for unlearning makes it vulnerable to adversarial prompts.
We addressed the vulnerability of existing unlearning T2I methods to prompt-based adversarial attacks. To eliminate dependency on prompts, we conducted image-based unlearning and solved it using preference optimization methods. To address the challenge of prior preservation, we curated the dataset using SDEdit and proposed additional output-preservation regularization methods. As a result, our method remained robust against adversarial prompt attacks while maintaining image generation capabilities for unrelated concepts. Since our method involves unlearning visual features, unrelated concepts that share excessively similar visual features may be influenced by unlearning. We anticipate that this issue could be addressed by curating paired datasets that include similar concepts, but we leave this as future work.
Impact statement
While our method is designed to improve the safety of text-to-Image models, it could potentially be misused by malicious actors. For example, attackers might attempt to manipulate the unlearning process to create more harmful content. Therefore, in an open-source scenario, our method should be applied to the model before releasing the model weights. Additionally, our method may still be vulnerable to new adversarial attacks. Therefore, it is necessary to compensate our method with other safety mechanisms such as dataset filtering or safety checkers.
Appendix A Ablation study
KL-constrained optimization
In this section, we study the effect of KL-constrained optimization. When $\beta\rightarrow 0$, the sigmoid function can be approximated as a linear function, then DUO approaches a naïve combination of gradient ascent for preferred samples and gradient descent for de-preferred samples [65]. As shown in Figure 15, when $\beta\rightarrow 0$, the images are somewhat degraded, and the generation results for unrelated concepts also change. On the other hand, when KL divergence regularization is present, we can confirm that the generation results for unrelated concepts hardly change at all.
Figure 15: Ablation study on KL divergence regularization.
Appendix B Implementation details
B.1 Dataset generation
Nudity
For generating the unsafe images $x^{-}$, we use “naked” as prompt with a guidance strength of 7.5. For $x^{+}$, we use “dressed” as with a strength of 7.5, and use “naked” as a negative prompt. We generate 256 images. We use nudenet [2] to regenerate images if there was no nudity in $x^{-}$ or if there was nudity in $x^{+}$.
Violence
To generate violence images we use four sub-concepts: Blood, Suffering, Gun, and Horror. To generate unsafe images $x^{-}$, we use the prompt: “a man, bleeding”, “a man, suffering”, “a man, with a gun”, “horror”. We generate 64 image per each prompt, total 256 images. For safe images $x^{+}$, we use the prompt: “a man, in comfort”, “a man, enjoying”, “a man, with peace”, “comfort”, with using unsafe prompt as a negative prompt.
SDEdit
When we use SDEdit, the magnitude of the added noise is $t=0.75T$, where T is the maximum diffusion timesteps, and the guidance scale used is 7.5.
B.2 Optimization.
DUO
Since $\beta$ linearly multiplies the learning rate at the time near initialization, we rescale the learning rate, dividing it by the same amount each time we increase $\beta$. Using $\beta=100$ as a baseline, we use a learning rate of 1e-4 for nudity and 2e-5 for violence. We use a LoRA rank of 32 for nudity and 2048 for violence. When conducting violence unlearning, we generally observe that the training is more unstable compare to nudity, and we mitigate this by increasing the LoRA rank and lowering the learning rate. Exploring the relationship between dataset characteristics and optimization stability in DUO is an interesting avenue for future work.
Concept Inversion
For Textual Inversion training, we used the Adam optimizer with a learning rate of lr=5.0e-03, batch size 4, and 3000 gradient steps. We used the same hyperparameters for all unlearning models.
Appendix C Derivation
Eq. (10) $\rightarrow$ Eq. (17)
Start from Eq. (10).
$L_{\text{Diffusion-DPO}}=-\mathbb{E}_{x_{0}^{+},x_{0}^{-}}[\log\sigma{(}\mathbb{E}_{p_{\theta}(x_{1:T}^{+}|x_{0}^{+}),p_{\theta}(x_{1:T}^{-}|x_{0}^{-})}[\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{+})}{p_{\phi}(x_{0:T}^{+})}}-\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{-})}{p_{\phi}(x_{0:T}^{-})}}]{)}]$ (14)
Since $-\log\sigma(\cdot)$ is a convex function, we can utilize Jensen’s inequality.
$L_{\text{Diffusion-DPO}}$ $\leq-\mathbb{E}_{x_{0}^{+},x_{0}^{-},p_{\theta}(x_{1:T}^{+}|x_{0}^{+}),p_{\theta}(x_{1:T}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{+})}{p_{\phi}(x_{0:T}^{+})}}-\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{-})}{p_{\phi}(x_{0:T}^{-})}}]{)}\right]$ (15a)
$=-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{1:T}^{+}|x_{0}^{+}),q(x_{1:T}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{+})}{p_{\phi}(x_{0:T}^{+})}}-\beta\log{\frac{p_{\theta}^{*}(x_{0:T}^{-})}{p_{\phi}(x_{0:T}^{-})}}]{)}\right]$ (15b)
In (15b), the intractable $p_{\theta}(x_{1:T}^{+}|x_{0}^{+})$ and $p_{\theta}(x_{1:T}^{-}|x_{0}^{-})$ are approximated by the diffusion forward process $q(x_{1:T}|x_{0})$. Now, let’s apply the product rule, i.e., $p(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p(x_{t-1}\mid x_{t})$, to $p_{\theta}(x_{0:T})$ and $p_{\phi}(x_{0:T})$.
$L_{\text{Diffusion-DPO}}$ $\leq-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{1:T}^{+}|x_{0}^{+}),q(x_{1:T}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\log{\frac{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}{p_{\phi}(x_{t-1}^{+}|x_{t})}}-\log{\frac{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}{p_{\phi}(x_{t-1}^{-}|x_{t})}}]{)}\right]$ (16a)
$=-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{1:T}^{+}|x_{0}^{+}),q(x_{1:T}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}(\log{\frac{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}{q(x_{t-1}^{+}|x_{t})}}-\log{\frac{p_{\phi}^{*}(x_{t-1}^{+}|x_{t})}{q(x_{t-1}^{+}|x_{t})}})-(\log{\frac{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}{q(x_{t-1}^{-}|x_{t})}}-\log{\frac{p_{\phi}^{*}(x_{t-1}^{-}|x_{t})}{q(x_{t-1}^{-}|x_{t})}})]{)}\right]$ (16b)
$=-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{t}^{+}|x_{0}^{+}),q(x_{t}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\bigg{(}\begin{aligned} &\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{+}|x_{t})\big{)}-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{+}|x_{t})\big{)}\\ &-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{-}|x_{t})\big{)}-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{-}|x_{t})\big{)}\end{aligned}\bigg{)}]{)}\right]$
Finally, by applying the variational lower bound of the diffusion loss, we obtain the upper bound of the above equation as follows [23]:
$L_{\text{Diffusion-DPO}}\leq$ $-\mathbb{E}_{(x_{t}^{+},x_{t}^{-})\sim D,x_{t}^{+}\sim q(x_{t}^{+}|x_{0}^{+}),x_{0}^{-}\sim q(x_{t}^{-}|x_{0}^{-})}$ (17)
$\left[\log\sigma\left(-\beta T\omega(\lambda_{t})\left(\begin{aligned} &\|\epsilon-\epsilon_{\theta}(x_{t}^{+},t)\|_{2}^{2}-\|\epsilon-\epsilon_{\phi}(x_{t}^{+},t)\|_{2}^{2}\\ &-\left(\|\epsilon-\epsilon_{\theta}(x_{t}^{-},t)\|_{2}^{2}+\|\epsilon-\epsilon_{\phi}(x_{t}^{-},t)\|_{2}^{2}\right)\end{aligned}\right)\right)\right]$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: As we claimed in the abstract, we experimentally verify our unlearning method in §$\S$ 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: §$\S$ 5 discusses the limitations of image-based unlearning.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: In this paper, there are no theory to prove. This is experimental paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In the Appendix B, we provide every details to reproduce our results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will publicly open the source-code for reproducible.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe all the hyperparameters, optimizer, and dataset in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to the lack of computational resources, it is infeasible to compute the error bar by taking an experiment on multiple random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report computation resources and time complexity of our algorithm in Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: In Appendix 5, we address potential risk of this work.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper involves impact statements that unlearning research can be abused by malicious users. Our research should be applied to the model so that it communicates with users only through API communication or before the open-source model is deployed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: In our paper, we plan to release a model with not-safe-for-work (NSFW) concepts removed, so there is no high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The Stable Diffusion, I2P benchmark, Sneakprompt, Ring-A-Bell, and Concept Inversion that we used are all open source.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We haven’t released the code yet. When we release it later, we will provide proper documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: In this study, we do not conduct a user study.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: In this study, we do not conduct a user study.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.